OpenStudy (anonymous):

Given the triangle pictured, use the Pythagorean Theorem to find the missing side. Round your solution to the nearest tenth, if necessary.

OpenStudy (anonymous):



OpenStudy (anonymous):

A. c = 10 B. c = 8.5 approximately C. c = 12 D. c = 11.4 approximately

OpenStudy (anonymous):

a^2 + b^2 = c^2 8^2 + 9^2 = c^2 64 + 81 = c^2 145 = c^2 sqrt(145) = c 12.0 = c

OpenStudy (anonymous):

so whic one would it be?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

aha you just want the answer? w

OpenStudy (anonymous):

no, i didnt understand it

OpenStudy (anonymous):

pythagorean thereom is a^2 + b^2 = c^2 a and b are the two legs of the triangle and c is the hypotenuse, in the thing that you drew; a= 8, b =9, and c = c or a = 9, b = 8. so just plug and chug to solve for c.

OpenStudy (anonymous):

okay thanks very much:)

OpenStudy (anonymous):

yeap no problem